# MetaToken: Detecting Hallucination in Image Descriptions by Meta Classification

### v1

@@ 5.3 Metric Analysis
In this section, we investigate the information contained in our proposed metrics introduced in Sec. 3.2. A visualization of these metrics can be found in Appendix C. We make use of the least absolute shrinkage and selection operator (LASSO) algorithm [12, 58] to analyze the predictive power of the metrics considered. The LASSO method performs a variable selection for a linear regression including the estimation of the corresponding coefficients ranking the most informative metrics. For the attention metrics (Eq. 5), we use the maximum of the absolute values of all $G$ weight coefficients. The LASSO path based on the image captions obtained from mPLUG-Owl using nucleus sampling is plotted in Fig. 3. The attention metrics $A^{g},g=0,\dots,G-1$ (Eq. 5) are selected first, closely followed by the absolute occurrence $N$ (Eq. 4), the log probability $L$ (Eq. 6) as well as the cumulated log probability $C$ (Eq. 7). Moreover, the LASSO path indicates a minor relevance of the co-occurrence score $H$ (Eq. 2), the sequence score $S$ (Eq. 8) and the variance $V$ ((9)). We obtain similar results independently from the underlying LVLM or generation method (see Fig. 5 in Appendix B). Tab. 7 lists the average rank of all metrics contained in the metric set $\mathcal{M}$ (Eq. 15) during the LASSO selection. While most of the metrics are selected during the LASSO path indicated through non-zero coefficients (see Fig. 5), Fig. 4 shows that five metrics are usually enough to achieve high $AUPRC$ values. Further metrics only add minor additional information to the classifier. Moreover, we provide the LASSO path including the CLIPScore (Eq. 14) for mPLUG-Owl [67] and LLaVa [23] in Appendix B. While for mPLUG-Owl, the CLIPScore is selected first, our proposed absolute occurrence $N$ (Eq. 4) is the first metric selected in the LLaVa path underlining the informative content induced by our metric.
@@
Finally, we refer to Appendix C to emphasise the importance of our statistical analysis based on the LASSO algorithm. While the relative position $P$ (Eq. 3) and the probability margin $M$ (Eq. 12) might look like a proper metric to classify between hallucinated and true objects (see Fig. 7), our analysis shows that these metrics only add minor information to the classifier reflected by an average rank of $10.25$ and $11.50$, respectively (see Tab. 7).
@@ 5.4 MetaToken and Revision of Image Descriptions
Table 8: Integration of MetaToken into LURE. Results of MetaToken (Ours) plugged into the LURE mitigation method [72]. The superscripts denote the hallucination recall values for the respective method. PR and FPR denote the hallucination precision and false positive rate, respectively. The best results are highlighted, the second best results are underlined.
@@
In this section, we investigate MetaToken as a substitute for the LURE detection for InstructBLIP [8] and mPLUG-Owl [67]. LURE [72] serves as a hallucination mitigation method using a MiniGPT-4-based revisor to rectify image captions. To this end, LURE applies thresholds on the log probability $L$ (Eq. 6) and the relative position $P$ (Eq. 3) to detect possible object hallucinations and replace them by the I-don’t-know string "[IDK]". The resulting image caption and the input image are fed into the revisor afterwards to rectify the detected tokens. In our experiments, we replace the threshold-based LURE detection with our proposed MetaToken method (see Sec. 3). Since LURE evaluates all objects instead of being restricted to the MSCOCO dataset [37], we train our classifier on the general set of metrics
=$\begin{split}\mathcal{M}^{*}=\{P_{o_{j}},N_{o_{j}},A^{0}_{o_{j}},\dots,A^{G-1}_{o_{j}},L_{o_{j}},\\ C_{o_{j}},S_{o_{j}},V_{o_{j}},E_{o_{j}},R_{o_{j}},M_{o_{j}},D_{o_{j}}\}\end{split}$ (23)
abstaining from the MSCOCO-related metrics $I$ (Eq. 1) and $H$ (Eq. 2). The results are stated in Appendix D. In accordance with Fig. 4, we observe almost no performance drop when considering the general metric set $\mathcal{M}^{*}$.
@@
First of all note that as in [70], we are not able to reproduce the results in [72] with respect to InstructBLIP. While we achieve a hallucination rate of $10.4\%$ for InstructBLIP captions (see Tab. 3), the rectified image captions by LURE include $10.9\%$ of hallucinations, even increasing the amount of hallucinated objects (see Tab. 8). We believe that this observation results from the fact that MiniGPT-4 has a higher hallucination rate than InstructBLIP (see Tab. 3). Since the LURE detection has a false positive rate (FPR) of $42.2\%$, as stated in Tab. 8, we guess that the MiniGPT-4-based revisor replaces true objects by hallucinated objects, even though the revisor is fine-tuned to mitigate hallucinations.
However, using our detection method, we are able to mitigate hallucinations achieving $\textrm{CHAIR}_{i}$ values of $9.8\%$. To this end, note that we can control the precision-recall-ratio in our method by varying the decision threshold of our lightweight meta classifier. For a recall of $70\%$, we observe a FPR of $9.4\%$ only. Thus, we prevent the revisor from including additional hallucinations by replacing false positives, that is, true objects. The results in Tab. 8 confirm our assumption on InstructBLIP: The higher the FPR, the higher the number of hallucinations induced by the revisor.
@@
On the other hand, LURE reduced the number of hallucinations for mPLUG-Owl by $52.98\%$, i.e. from $30.2\%$ to $14.2\%$ with a FPR of $46.4\%$. Since the revisor induces substantially less hallucinations than mPLUG-Owl, the correction of true positives outweights the potential introduction of new hallucinations by replacing false positives. In fact, we observe from Tab. 8 that higher recall values, and thus, higher FPRs, lead to consistently lower hallucination rates for mPLUG-Owl. However, note that the superior precision-recall-ratio from our approach also outperforms the LURE results on mPLUG-Owl: For a recall of $80\%$ (which is closest to the LURE detection of $78.6\%$), we reduce the proportion of hallucinations by $56.62\%$ to $13.1\%$, that is, $3.64\%$ less hallucinations compared to the LURE baseline.
@@ 6 Limitations
While MetaToken achieves high classification performance, it still has room for improvement for strong LVLMs inducing only few hallucinations. For these models, a high hallucination recall rate can be achieved at the cost of a high false positive rate. Moreover, our method is restricted to the problem of object instance hallucination detection in image captions, while the detection of attribute-level hallucinations still remains an unsolved problem we will tackle in future work. Lastly, our method is based on the handcrafted CHAIR method [52] based on human annotations. While this evaluation method relies on exact matching, it still leads to mismatches due to misinterpretations of the generated language output of LVLMs.
@@ 7 Conclusion
In this paper, we introduce MetaToken, a novel lightweight hallucination detection technique for LVLMs based on meta classification. Inspired by recently discovered causes of hallucinations, we propose and analyse a broad set of potential factors for hallucinations in LVLMs. Based on a comprehensive statistical analysis of these factors, we reveale key indicators of hallucinations which have been overseen in previous works. We evaluate our method on four state-of-the-art LVLMs and different generation configurations achieving $AUROC$ values of up to $92.19\%$ and $AUPRC$ values of up to $82.69\%$. Moreover, we show that our lightweight classifier reliably detects hallucinations inducing an $ECE$ between $0.81\%$ and $2.01\%$. Finally, we demonstrate that MetaToken can be easily integrated into the LURE mitigation method reducing the hallucination rate by up to $56.62\%$, i.e., $3.64\%$ less hallucinations than the LURE baseline. As future work, we will tackle the problem of attribute-level hallucination detection for general visual question answering tasks.
@@
Supplementary Material
Appendix A Meta Classifier Configurations
In our experiments, we use the sklearn.linear_model.LogisticRegression [^2] https://scikit-learn.org/1.3/modules/generated/sklearn.linear_model.LogisticRegression.html and sklearn.ensemble.GradientBoostingClassifier [^3] https://scikit-learn.org/1.3/modules/generated/sklearn.ensemble.GradientBoostingClassifier.html methods with scikit-learn version $1.3.2$. Tab. 9 states the applied configurations. For those parameters not listed in Tab. 9, we refer to the default settings.
Appendix B LASSO Path
(a) InstructBLIP (sample)
(b) mPLUG-Owl (sample)
(c) InstructBLIP (beam)
(d) mPLUG-Owl (beam)
(e) MiniGPT-4 (sample)
(f) LLaVa (sample)
Figure 5: LASSO path. LASSO path for the set of metrics $\mathcal{M}$ (Eq. 15). $A$ denotes the maximum of the absolute values of all $G$ weight coefficients for the attention metrics $A^{g},g=0,\dots,G-1$ (Eq. 5). (beam) and (sample) reflect the beam search and nucleus sampling algorithm, respectively.
(a) mPLUG-Owl (sample)
(b) LLaVa (sample)
Figure 6: LASSO path. LASSO path for the set of metrics $\mathcal{M}^{\mathrm{clip}}$ (Eq. 16). $A$ denotes the maximum of the absolute values of all $G$ weight coefficients for the attention metrics $A^{g},g=0,\dots,G-1$ (Eq. 5). (beam) and (sample) reflect the beam search and nucleus sampling algorithm, respectively.
We analyze the informative content included in the metrics proposed in Sec. 3.2 for the considered LVLMs InstructBLIP (Vicuna-7B) [8], mPLUG-Owl (LLaMA-7B) [67], MiniGPT-4 (Vicuna-7B) [73], and LLaVa 1.5 (Vicuna-7B) [23] based on nucleus sampling [22]. For InstructBLIP and mPLUG-Owl we further investigated the considered metrics based on beam search [3]. To this end, we make use of the least absolute shrinkage and selection operator (LASSO) algorithm [12, 58] to analyze the predictive power of the metrics considered. The LASSO method performs a variable selection for a linear regression including the estimation of the corresponding coefficients ranking the most informative metrics. Fig. 5 shows the LASSO paths for the considered LVLMs and generation configurations. While the metrics absolute occurrence $N$ (Eq. 4), mean absolute attention $A$ (Eq. 5), log probability $L$ (Eq. 6) and cumulated log probability $C$ (Eq. 7) are selected first in almost all settings, we observe different metric ranks between the sampling-based captions and beam search-based captions. While the coefficient for the entropy $E$ (Eq. 10) is vanishing for the sampled captions (see Figs. 5(a), 5(b), 5(e) and 5(f)), it is selected as the sixth metric for the beam search-based captions (see Figs. 5(c) and 5(d)).
Table 9: Classifier Configurations. The configurations applied in our experiments for the sklearn.linear_model.LogisticRegression (LR) and sklearn.ensemble.GradientBoostingClassifier (GB) classifier.
LR GB
scikit-learn $1.3.2$ $1.3.2$
random_state 00 00
solver saga -
max_iter $1000$
tol $31\mathrm{e}-$ 3 -
Moreover, we investigate the LASSO paths for the extended metric set $\mathcal{M}^{\mathrm{clip}}$ (Eq. 16) for the LVLMs sharing the vision encoder with CLIP [50], that is, mPLUG-Owl and LLaVa in Fig. 6. For mPLUG-Owl, the $CLIP$ metric (Eq. 14) is selected as the first metric as we can see in Fig. 6(a) followed by the mean absolute attention $A$ (Eq. 5) the absolute occurrence $N$ (Eq. 4) and the log probability $L$ (Eq. 6). For LLaVa, on the other hand, our proposed metric absolute occurrence $N$ (Eq. 4) is selected first, followed by the $CLIP$ metric (Eq. 14), the log probability $L$ (Eq. 6) and cumulated log probability $C$ (Eq. 7).
Appendix C Metric Visualization
In order to provide a visual analysis of our proposed metrics (see Sec. 3.2), we plot our metrics for true and hallucinated objects in Figs. 7 and 8. While some of the attention metrics $A^{g},g=0,...,31$ fully overlap for true and hallucinated objects (see Fig. 8), for specific attention heads like $g=9,24$ and $28$ the respective metrics distinguish between hallucinated and true objects. We can also see in Fig. 7 that our proposed metric absolute occurrence $N$ (Eq. 4) has a high density for $N=1$ for hallucinated objects, that is, if the object is only mentioned once in the image caption. True objects usually occur several times in the image caption. However, we want to emphasise the importance of our statistical analysis based on the LASSO algorithm in Appendix B. While in Fig. 7, the relative position $P$ (Eq. 3) might look like a better metric to classify between hallucinated and true objects than the log probability $L$ (Eq. 6) or the cumulated log probability $C$ (Eq. 7), our analysis in the previous section shows that the relative position $P$ only add minor information to the classifier reflected by a rather small coefficient compared to the log probability $L$ and cumulated log probability $C$ (see Fig. 5).
Figure 7: Metrics. Visualization of a selection of our proposed input metrics defined in Sec. 3.2 for true and hallucinated objects.
Figure 8: Attention Metrics. Visualization of the attention metrics Eq. 5 defined in Sec. 3.2 for true and hallucinated objects.
Appendix D Metric Combination
Since recent works [36, 72] have shown that LVLMs are prone to hallucinate (co-occurring) MSCOCO objects [37], we include the MSCOCO class index $I$ (Eq. 1) and the co-occurrence score $H$ (Eq. 2) into our proposed set of metrics $\mathcal{M}$. However, in order to provide a general method, which can be applied to any dataset without any restrictions, we investigate the performance of MetaToken with respect to the general set of metrics $\mathcal{M}^{*}$ (Eq. 23) abstaining from the MSCOCO-related metrics $I$ and $H$. The results are provided in Tab. 10. Even though our exploration from Appendix B shows that the MSCOCO class index $I$ is usually under the top five metrics selected, we observe almost no performance drop in the classification results for $\mathcal{M}^{*}$.
Table 10: Experimental Results for different Set of Metrics. Hallucination detection results for InstructBLIP [8] and mPLUG-Owl [67] for the MSCOCO [37] specific metric set $\mathcal{M}$ (Eq. 15) and the general set $\mathcal{M}^{*}$ (Eq. 23). The classifier based on logistic regression and gradient boosting are denoted by LR and GB, respectively. The best results are highlighted.
LR GB LR GB LR GB
InstructBLIP $\mathcal{M^{*}}$ $\textbf{91.34}^{(\pm 1.7\mathrm{e}-3)}$ $\textbf{91.49}^{(\pm 1.8\mathrm{e}-3)}$ $89.93^{(\pm 8.9\mathrm{e}-3)}$ $89.93^{(\pm 7.3\mathrm{e}-3)}$ $56.07^{(\pm 1.2\mathrm{e}-2)}$ $56.71^{(\pm 7.6\mathrm{e}-3)}$
$\mathcal{M}$ $91.33^{(\pm 1.8\mathrm{e}-3)}$ $91.48^{(\pm 1.5\mathrm{e}-3)}$ $\textbf{89.94}^{(\pm 9.1\mathrm{e}-3)}$ $\textbf{90.39}^{(\pm 6.8\mathrm{e}-3)}$ $\textbf{56.13}^{(\pm 1.2\mathrm{e}-2)}$ $\textbf{57.50}^{(\pm 1.0\mathrm{e}-2)}$
mPLUG-Owl $\mathcal{M^{*}}$ $\textbf{82.90}^{(\pm 1.9\mathrm{e}-3)}$ $83.26^{(\pm 2.6\mathrm{e}-3)}$ $88.41^{(\pm 3.9\mathrm{e}-3)}$ $88.9^{(\pm 2.8\mathrm{e}-3)}$ $75.94^{(\pm 6.2\mathrm{e}-3)}$ $77.04^{(\pm 5.8\mathrm{e}-3)}$
$\mathcal{M}$ $82.87^{(\pm 2.7\mathrm{e}-3)}$ $\textbf{83.56}^{(\pm 3.1\mathrm{e}-3)}$ $\textbf{88.55}^{(\pm 3.7\mathrm{e}-3)}$ $\textbf{89.87}^{(\pm 2.3\mathrm{e}-3)}$ $\textbf{76.11}^{(\pm 5.8\mathrm{e}-3)}$ $\textbf{78.61}^{(\pm 7.9\mathrm{e}-3)}$

### v2

@@ 5.3 Feature Analysis
In this section, we investigate the information contained in our proposed input features introduced in Sec. 3.2. A visualization of these statistics can be found in Sec. 8.5. We make use of the least absolute shrinkage and selection operator (LASSO) algorithm [Efron et al., 2004, Tibshirani, 2018] to analyze the predictive power of the input features considered. The LASSO method performs a variable selection for a linear regression including the estimation of the corresponding coefficients ranking the most informative features. For the attention features (Eq. 3), we use the maximum of the absolute values of all $G$ weight coefficients. Fig. 3 shows the LASSO path for mPLUG-Owl. Our proposed attention features $A^{g},g=0,\dots,G-1$ (Eq. 3) are selected first, closely followed by the absolute occurrence $N$ (Eq. 2), the log probability $L$ (Eq. 4) as well as the cumulated log probability $C$ (Eq. 5). Moreover, the LASSO path indicates a minor relevance of the sequence score $S$ (Eq. 6) and the variance $V$ (Eq. 7) indicated through vanishing coefficients. We obtain similar results independently from the underlying LVLM or dataset (see Sec. 8.4). Tab. 7 lists the average rank of all features contained in the feature set $\mathcal{M}$ (Eq. 12) during the LASSO selection. While most of the features are selected during the LASSO paths indicated through non-zero coefficients (see Fig. 5), Fig. 4 shows that four features are usually enough to achieve high $AUPRC$ values. Further features only add minor additional information to the classifier.
@@
Finally, we refer to Sec. 8.5 to emphasize the importance of our statistical analysis based on the LASSO algorithm. While the relative position $P$ (Eq. 1) and the probability margin $M$ (Eq. 10) might look like proper features to classify between hallucinated and true objects (see Fig. 8), our analysis shows that these features only add minor information to the classifier reflected by an average rank of $8.750$ and $8.375$, respectively (see Tab. 7).
Figure 4: $AUPRC$ as a Function of the Number of Features. The classification performance of MetaToken in terms of $AUPRC$ as a function of the number of features for different LVLMs. The features are selected along the LASSO path of the respective LVLM.
@@ 5.4 MetaToken and Revision of Image Descriptions
Table 8: Integration of MetaToken into LURE. Results of MetaToken (Ours) plugged into the LURE mitigation method [Zhou et al., 2023]. The superscripts denote the hallucination recall values for the respective method. PR and FPR denote the hallucination precision and false positive rate, respectively. The best results are highlighted, the second best results are underlined.
@@
In this section, we investigate MetaToken as a substitute for the LURE detection on the MSCOCO dataset. LURE [Zhou et al., 2023] serves as a hallucination mitigation method using a MiniGPT-4-based revisor to rectify image captions. To this end, LURE applies thresholds on the log probability $L$ (Eq. 4) and the relative position $P$ (Eq. 1) to detect possible object hallucinations and replace them by the I-don’t-know string "[IDK]". The resulting image caption and the input image are fed into the revisor afterwards to rectify the detected tokens. In our experiments, we replace the threshold-based LURE detection with our proposed MetaToken method (see Sec. 3).
@@
First of all note that as in [Zhao et al., 2024], we are not able to reproduce the results in [Zhou et al., 2023] with respect to InstructBLIP. While we achieve a hallucination rate of $10.4\%$ for InstructBLIP captions (see Tab. 2), the rectified image captions by LURE include $10.9\%$ of hallucinations, even increasing the amount of hallucinated objects (see Tab. 8). We believe that this observation results from the fact that MiniGPT-4 has a higher hallucination rate than InstructBLIP (see Tab. 2). Since the LURE detection has a false positive rate (FPR) of $42.2\%$ (see Tab. 8), i.e., detects true objects as hallucinations, we guess that the MiniGPT-4-based revisor replaces true objects by hallucinated objects, even though the revisor is fine-tuned to mitigate hallucinations. However, applying our detection method, we are able to mitigate hallucinations achieving $\textrm{CHAIR}_{i}$ values of $9.8\%$. To this end, note that we can control the precision-recall-ratio in our method by varying the decision threshold of our lightweight meta classifier. For a recall of $70\%$, we observe a FPR of $9.4\%$ only. Thus, we prevent the revisor from including additional hallucinations by replacing false positives, that is, true objects. The results in Tab. 8 confirm our assumption on InstructBLIP: The higher the FPR, the higher the number of hallucinations induced by the revisor.
@@
For mPLUG-Owl, LURE reduced the number of hallucinations by $52.98\%$, i.e. from $30.2\%$ to $14.2\%$ with a FPR of $46.4\%$. Since the revisor induces substantially less hallucinations than mPLUG-Owl, the correction of true positives outweighs the potential introduction of new hallucinations by replacing false positives. In fact, we observe from Tab. 8 that higher recall values, and thus, higher FPRs, lead to consistently lower hallucination rates for mPLUG-Owl. However, note that the superior precision-recall-ratio from our approach, again, outperforms the LURE results: For a recall of $80\%$ (which is closest to the LURE detection of $78.6\%$), we reduce the proportion of hallucinations by $56.62\%$ to $13.1\%$, that is, $3.64\%$ less hallucinations compared to the LURE baseline.
@@ 6 LIMITATIONS
Due to the lack of automated and reproducible token-level evaluation methods for attribute-level hallucinations, MetaToken is currently restricted to the problem of object hallucination detection, while the detection of attribute-level hallucinations remains an unsolved problem we will tackle in future work. Moreover, while the automated CHAIR method [Rohrbach et al., 2018] relies on ground truth labels, it still leads to mismatches due to misinterpretations of the generated language output of LVLMs.
@@ 7 CONCLUSION
In this paper, we introduce MetaToken, a novel lightweight hallucination detection technique for LVLMs based on meta classification. Inspired by recently discovered causes of hallucinations, we propose and analyze a broad set of potential factors for hallucinations in LVLMs. Based on a comprehensive statistical analysis of these factors, we reveal key indicators of hallucinations. We evaluate our method on four SOTA LVLMs achieving $AUROC$ values of up to $92.12\%$ and $AUPRC$ values of up to $84.01\%$. Moreover, we show that our lightweight classifier detects hallucinations inducing an $ECE$ between $0.81\%$ and $2.01\%$. Finally, we demonstrate that MetaToken can be easily integrated into the LURE mitigation method reducing the hallucination rate by up to $56.62\%$, i.e., $3.64\%$ less hallucinations than the LURE baseline. As future work, we will tackle the problem of attribute-level hallucination detection for general visual question answering tasks.